The heights of five famous human-made structures are related as follows:
•The height of the Statue of Liberty is 65 feet more than half the height of the Great Pyramid at Giza.
•The height of the Eiffel Tower is 36 feet more than three times the height of Big Ben.
•The Great Pyramid at Giza is 164 feet taller than Big Ben.
•The Leaning Tower of Pisa is 137 feet shorter than Big Ben.
•The total of all the heights is nearly half a mile. In fact, the sum of the five heights is 2264 feet.
Find the height of each of the five structures.

Answer by scott8148(6628)

let x=Ben, so Eiffle=3x+36, Giza=x+164, Pisa=x-137, and Liberty=.5(x+164)+65

x+(3x+36)+(x+164)+(x-137)+(.5x+82+65)=2264 __ 6.5x+210=2264 __ 6.5x=2054